A positive test charge of 5.0 × 10–4 C is in an electric field that exerts a force of 2.5 × 10–4 N on it. What is the magnitude of the electric field at the location of the test charge?
1
Expert's answer
2011-08-25T13:12:01-0400
We have that F=Eq, whence
E = F / q = 2.5 × 10–4 N / 5.0 × 10–4 C = 0.5 N/C